Suppose your firm is considering two mutually exclusive, required projects with the cash flows shown below. The required rate of return on projects of both of their risk class is 8 percent, and that the maximum allowable payback and discounted payback statistic for the projects are 2 and 3 years, respectively. Time: 0 1 2 3 Project A Cash Flow -30,000 20,000 40,000 11,000 Project B Cash Flow -40,000 20,000 30,000 60,000 Use the NPV decision rule to evaluate these projects; which one(s) should be accepted or rejected? rev: 12_04_2012  accept A, reject B  accept both A and B  accept neither A nor B  reject A, accept B
Answer #1

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

A:

Present value of inflows=20000/1.08+40000/1.08^2+11000/1.08^3

=$61544.23

NPV=Present value of inflows-Present value of outflows

=(61544.23-30000)=$31544.23(Approx)

B:

Present value of inflows=20000/1.08+30000/1.08^2+60000/1.08^3

=$91868.62

NPV=$91868.62-$40000

=$51868.62

Hence since projects are mutually exclusive;B must be chosen and A rejected having higher NPV.